### v4

# The Generalization Ridge: Information Flow in Natural Language Generation

Abstract
Transformer-based language models have achieved state-of-the-art performance in natural language generation (NLG), yet their internal mechanisms for synthesizing task-relevant information remain insufficiently understood. While prior studies suggest that intermediate layers often yield more generalizable representations than final layers, how this generalization ability emerges and propagates across layers during training remains unclear. To address this gap, we propose InfoRidge, an information-theoretic framework, to characterize how predictive information—the mutual information between hidden representations and target outputs—varies across depth during training. Our experiments across various models and datasets reveal a consistent non-monotonic trend: predictive information peaks in intermediate layers—forming a generalization ridge—before declining in final layers, reflecting a transition between generalization and memorization. To further investigate this phenomenon, we conduct a set of complementary analyses that leverage residual scaling, attention pattern, and controlled model capacity to characterize layer-wise functional specialization. We further validate our findings with multiple-token generation experiments, verifying that the observed ridge phenomenon persists across decoding steps. Together, these findings offer new insights into the internal mechanisms of transformers and underscore the critical role of intermediate layers in supporting generalization.
1 Introduction
Transformer-based language models generate text by predicting tokens autoregressively, and they have achieved remarkable performance across a wide range of natural-language uses (Vaswani et al., 2017; Dong et al., 2022). Nevertheless, we lack a rigorous understanding of how these models acquire and synthesize task-relevant information during training.
A growing body of research has shown that intermediate layers in deep neural networks often surpass final layers in terms of representational quality and generalization performance (Liu et al., 2019b; Voita et al., 2019; Ansuini et al., 2019; Ahrens et al., 2023; Uselis and Oh, 2025). In language models, intermediate layers often encode richer semantic and more robust features than final layers (Fan et al., 2024; Jin et al., 2024; Skean et al., 2025). However, questions still remain: In NLG, how does information evolve across layers during training, and how are different layers of the network functionally organized to support generalization versus memorization?
To investigate these questions, we propose InfoRidge, an information-theoretic framework, to analyze information flow in language models. Building on matrix-based mutual information estimation (Giraldo et al., 2014), our approach quantifies how predictive signals transform across layers during training. We center our analysis on two complementary quantities: predictive information, defined as the mutual information $I(Z_{\ell};Y)$ between the hidden representation $Z_{\ell}$ at layer $\ell$ and the next-token label $Y$, reflecting how much task-relevant information is preserved; incremental information gain, denoted as $I(\Delta Z_{\ell};Y)$, where $\Delta Z_{\ell}$ is the residual changes between successive layer $\ell$ and $\ell-1$, measuring the additional predictive information introduced by each transformer block.
Using InfoRidge, we uncover a non-monotonic trend: predictive information rises through the early and middle layers, peaks in the upper-middle layers, and then declines in the later layers. We name this phenomenon the generalization ridge, where the model encodes the more generalizable task-relevant information. This ridge marks a structural division of labor: intermediate layers concentrate generalizable features that transfer across distributions, while later layers increasingly specialize in task-specific memorization. Incremental information gain further shows that ridge layers introduce the largest increases in predictive information, marking them as key contributors to the emergence of the ridge. This analysis directly connects the information peak to generalizable behavior and clarifies how generalization and memorization are distributed across depth.
We further extend InfoRidge to multi-token natural language generation. By treating the full output sequence as a structured target and measuring mutual information between hidden states and the long-form target distribution, we show that the generalization ridge persists across decoding steps. Despite increasing mutual information as generation progresses, the layer-wise structure remains stable: predictive information consistently peaks in early-to-intermediate layers and declines in later layers.
To understand the formation of the generalization ridge, we analyze attention patterns, model capacity and residual scaling coefficients. Attention analysis shows ridge layers attend to tokens that capture broadly useful features, aligning with the information peak. We also investigate the conditions under which the ridge emerges. Our depth ablation results show that the ridge only emerges beyond a certain depth threshold. Below this threshold, predictive information increases monotonically, indicating that sufficient capacity is a prerequisite for generalization ridge to emerge. We also introduce residual scaling coefficients $\beta_{\ell}$—learnable scalars applied to each residual block with all other weights frozen—as functional probes of layer-wise reliance (Liu et al., 2019a; Menghani et al., 2024). When evaluated under distribution shift, models have lower reliance on the final layer and shifts importance toward the intermediate layers.
Contributions.
Our work provides a unified perspective on how predictive information is structured across depth in transformer-based language models for natural language generation tasks:
1. Our work tracks the evolution of predictive information throughout training, establishing a clear connection between predictive information flow and generalization. It reveals a non-monotonic peak in the middle layers, which we refer to as the generalization ridge. This pattern reflects a meaningful transition in representational focus and aligns with stronger generalization behavior.
2. We introduce InfoRidge, an information-theoretic framework that applies matrix-based mutual information estimation to autoregressive language models to analyze information flow.
3. We analyze attention patterns, model capacity and residual scaling coefficients, providing causal evidences for the generalization ridge phenomenon.
2 Related Work
Understanding how information is encoded and transformed across layers has been studied through probing classifiers (Alain and Bengio, 2016), attention flow (Vig and Belinkov, 2019), and information-theoretic approaches such as the information bottleneck (Shwartz-Ziv and Tishby, 2017), mutual information estimation (Goldfeld, 2019), and matrix-based entropy (Giraldo et al., 2014), offering different lenses to quantify representational capacity, abstraction, and invariance across layers.
A growing body of research has shown that intermediate layers in deep networks often outperform final layers in terms of representational quality and task performance (Ansuini et al., 2019; Yosinski et al., 2014; Uselis and Oh, 2025; Ahrens et al., 2023; Ando et al., 2023). In language models, mid-depth layers tend to capture richer semantic or robust features than output layers (Liu et al., 2019b; Voita et al., 2019; Jin et al., 2024; Fan et al., 2024). Transformer representations have long been observed to follow a structured progression from syntactic to semantic information, as shown by classical probing studies on linguistic knowledge and the reconstruction of the NLP pipeline (Liu et al., 2019b; Tenney et al., 2019). These findings challenge the assumption that deeper layers always yield better representations.
This pattern holds across settings such as transfer learning (Yosinski et al., 2014), continual learning (Ahrens et al., 2023), and out-of-distribution generalization (Uselis and Oh, 2025). Furthermore, recent work has evaluated representation quality using entropy, curvature, and invariance (Skean et al., 2025), while other studies have examined embedding drift and representational geometry (Merchant et al., 2020; Dar et al., 2022), analyzed memorization and factual recall (Haviv et al., 2022; Yu et al., 2023), and introduced causal perspectives on layer importance through mediation analysis and targeted interventions (Vig et al., 2020; Meng et al., 2022). Training-dynamics studies investigated how earlier models develop and refine semantic features across depth, providing an additional perspective on the evolution of layer-wise representations (Merchant et al., 2020; Kumar et al., 2023). Correlational probes (e.g., linear probes (Alain and Bengio, 2016)) measure only whether a feature can be decoded from a representation, which reflects correlation but not causal influence. In contrast, causal methods intervene on internal activations to test how changing a component alters the model’s prediction, thereby identifying true causal contribution rather than mere feature presence. Our residual-scaling approach aligns with this causal perspective at layer level.
However, the underlying causes and functional role of this phenomenon remain only partially understood, motivating further investigation. Our work addresses this gap by tracing the evolution of predictive information throughout training and establishing a clear connection between predictive information flow and generalization. We reveal a consistent non-monotonic peak in the middle layers—termed the generalization ridge—that reflects a meaningful transition in representational focus and aligns with stronger generalization behavior. Additionally, unlike prior work focused on classification tasks, we extend the analysis of information flow to generation tasks, enabling us to understand how generalization and memorization dynamics evolve for natural language generation setting from an information-theoretic perspective.
3 InfoRidge: Information Estimation Framework
Figure 1: Overview of InfoRidge. (1) Step 1: Extract internal representations at each layer and compute residual changes between successive layers. (2) Step 2: Extract the target token embedding. (3) Step 3: Using these representations, we estimate the layer-wise predictive information $I(Z_{\ell};Y)$ and the incremental information gain $I(\Delta Z_{\ell};Y)$, which jointly characterize how information evolves across depth.
We propose InfoRidge, an information-theoretic framework that uses mutual information to quantify how predictive information propagates through transformers layers during training in NLG.
Motivating Insight.
Prior work has shown that internal representations in deep neural networks tend to align most closely with the true data distribution at an intermediate layer (He et al., 2024). By employing the Wasserstein distance (Villani et al., 2008), this alignment is shown to reach a minimum at a specific depth—referred to as the generalization funnel layer. At this point, the Min Wasserstein Generalization Bound (He et al., 2024) ensures that the upper bound on the generalization gap—defined as the expected difference between the population and empirical risks—is minimized. This highlights the critical role of intermediate layers in supporting generalization.
Research Question.
Despite the insight from prior studies of deep neural networks in classification, it remains unclear whether this generalizes to Transformer-based language models in natural language generation (NLG). This motivates our central question: In NLG, how does information evolve across layers during training, and how are different layers of the network functionally organized to support generalization versus memorization?
Hypothesis.
Building on the insight, we hypothesize that there exists a specific intermediate transformer layer that encodes the most generalizable representations for next-token prediction, characterized by maximal mutual information with the target label.
Hypothesis: Generalization Ridge.  There exists an intermediate layer $\ell^{*}\in\{1,\dots,L\}$ such that the mutual information between the hidden state and the target label peaks at that layer:
$\ell^{*}=\arg\max_{\ell}I(Z_{\ell};Y).$
We refer to this layer as the generalization ridge. This ridge layer aligns most strongly with generalizable features and serve as robust predictors under distribution shift, whereas later layers increasingly specialize in memorization.
InfoRidge Overview.
To empirically investigate this hypothesis, we propose InfoRidge, an information estimation framework, that characterizes how predictive information evolves across transformer layers. Specifically, we estimate two key quantities:
  • Predictive Information $I(Z_{\ell};Y)$: the mutual information between the hidden state at layer $\ell$ and the target token. This quantity measures how much information about the true next token is contained in the layer’s full representation. A high value indicates that the layer encodes a strong and direct signal relevant to the prediction task.
  • Incremental Information Gain $I(\Delta Z_{\ell};Y)$: the information introduced by the residual transformation at layer $\ell$, where $\Delta Z_{\ell}=Z_{\ell}-Z_{\ell-1}$. This captures the additional predictive signal gained through the residual transformation at layer $\ell$, isolating how much new task-relevant information is introduced on top of the previous layer’s representation.
Together, these metrics allow us to track both the accumulation and transformation of task-relevant information throughout the network.
Notation and Setup.
We consider a transformer model with $L$ residual blocks. Given an input sequence $x_{1:T}$ of length $T$, $z^{(\ell)}_{i}\in\mathbb{R}^{d}$ denotes the hidden state at the last token position of the $i$th input sequence in layer $\ell$, for $\ell=1,\dots,L$. For next-token prediction, the ground-truth label is denoted by $y_{i}\in\mathbb{R}^{d}$, corresponding to the embedding of the true next token from the vocabulary $\mathcal{V}$. The residual transformation introduced at layer $\ell$ is defined as $\delta z_{i}^{(\ell)}=z_{i}^{(\ell)}-z_{i}^{(\ell-1)}$.
Across a batch of $N$ sequences, we collect the representations:
$Z_{\ell}=\{z^{(\ell)}_{i}\}_{i=1}^{N}\!,\;Y=\{y_{i}\}_{i=1}^{N}\!,\;\Delta Z_{\ell}=\{\delta z_{i}^{(\ell)}\}_{i=1}^{N},$
where all vectors are $\ell_{2}$-normalized.
Computational Flow Overview.
Figure 1 illustrates the workflow used to extract intermediate representations for information analysis. In the first forward pass, a batch of $N$ input sequences $x_{1:T}$ is fed into the transformer to obtain hidden states at each layer. From these, we extract the final-token representations $Z_{\ell}$ and compute the residual changes $\Delta Z_{\ell}$ by differencing consecutive layer outputs. In the second forward pass, each input is concatenated with its ground-truth next token $y$, and we extract the corresponding label embedding $Y$ from the output of the embedding layer. These representations are then used to compute two information-theoretic quantities: the Predictive Information $I(Z_{\ell};Y)$, and the Incremental Information Gain $I(\Delta Z_{\ell};Y)$. We estimate both $I(Z_{\ell};Y)$ and $I(\Delta Z_{\ell};Y)$ using Equation 1 and 2, detailed below.
Matrix-Based Mutual Information.
We apply the matrix-based framework (Giraldo et al., 2014) to estimate mutual information. Let $\mathcal{U}$ be a random variable, from which we draw a set of vectors $U=\{\mathbf{u}_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$. A positive-definite Gram matrix $G_{U}\in\mathbb{R}^{N\times N}$ is computed using a Gaussian kernel $\kappa$ with bandwidth set to 1 and the matrix is then trace-normalized to satisfy $\operatorname{tr}(G_{U})=1$. To assess robustness with respect to kernel selection, we additionally test the Laplacian and Polynomial kernels. While these kernels introduce quantitative shifts in the magnitude, the trend remains unchanged, see Appendix E. The matrix-based Rényi entropy (with order $\alpha=1$) is then given by:
$H(\mathcal{U})\approx H(U)=-\operatorname{tr}(G_{U}\log G_{U}).$ (1)
Specifically, $G_{U}$ is constructed with entries $(G_{U})_{ij}=\exp\left(-\frac{\|u_{i}-u_{j}\|^{2}}{2\sigma^{2}}\right)$ and then trace-normalized.
Let $G_{\mathcal{U}}$ and $G_{\mathcal{V}}$ be the trace-normalized Gram matrices for two random variables $\mathcal{U}$ and $\mathcal{V}$, respectively. The mutual information between them is computed as:
$I(\mathcal{U};\mathcal{V})$ $\approx I(G_{\mathcal{U}};G_{\mathcal{V}})$ (2)
$=H(G_{\mathcal{U}})+H(G_{\mathcal{V}})-H(G_{\mathcal{U}}\circ G_{\mathcal{V}}),$
where $\circ$” denotes the Hadamard product. Mathematical details are in Appendix A.
4 Experimental Setup
Models.
We evaluate four models: GPT–2 Small (117M) (Radford et al., 2019), GPT–2 Medium (345M) (Radford et al., 2019), Qwen–2.5 0.5B (Yang et al., 2024), and LLaMA 3.1 8B (Meta AI, 2024). All models are fine-tuned on NLG tasks. Each model shares weights between the input token embedding and the output language modeling head.
Datasets.
We assess model behavior across three tasks casted into NLG problems: CLUTRR (Sinha et al., 2019), a relational reasoning benchmark; ECQA (Aggarwal et al., 2021), a commonsense QA benchmark; CNN/DailyMail (See et al., 2017; Hermann et al., 2015), a summarization benchmark and Synthetic Arithmetic, a controlled dataset designed to disentangle task-relevant signal from noise. Dataset and implementation details are in Appendix B and C.
Synthetic Arithmetic Dataset Construction. We construct a synthetic dataset to separate signal learning from noise memorization. Each input is a sequence of 10 elements, where the signal follows an arithmetic progression modulo $K$, computed as $s_{t}=(s_{0}+t\cdot d)\bmod K$ with $s_{0}\in[0,K{-}1]$ and $d\in[1,K{-}1]$. Each element takes the form S{signal}_N{noise}, where noise is sampled from $\mathcal{U}_{\text{int}}(0,\texttt{noise\_range-1})$ ($\mathcal{U}_{\text{int}}$ denotes the uniform distribution). The model is trained to predict the signal value of the final (10th) element using the preceding elements as input context. For example, with $K=5$, $s_{0}=1$, and $d=2$, a sample input might be S1_N42 S3_N77...S2_N37, with the target signal being 4. By varying $K$, we can induce structured distribution shifts.
5 Generalization Ridge: Layer-wise Mutual Information Training Dynamics
Understanding how layers encode task-relevant information is key to uncovering the internal mechanisms that support generalization in deep language models. In this section, we trace the evolution of two complementary forms of mutual information—Predictive Information and Incremental Information Gain—across transformer depth and training time, revealing a consistent structure in information flow and highlighting the generalization–memorization trade-off.
5.1 Predictive Information: Information Peaks at Intermediate Layers
We investigate how predictive information—defined as the mutual information between hidden representations and target labels—evolves across the depth of transformer models. Specifically, for each layer $\ell$, we compute the matrix-based mutual information $I(Z_{\ell};Y)$ between the hidden state $Z_{\ell}$ and the next-token ground truth $Y$. This quantity measures how much task-relevant signal is retained in the representation as it propagates through the network. By tracing $I(Z_{\ell};Y)$ across layers, we obtain a layer-wise trajectory of information flow, which reveals not only where predictive content is preserved but also how it is transformed or diminished as the model processes.
(a) GPT-2 Small on Synthetic
(b) GPT-2 Medium on CLUTRR
(c) Qwen-2.5-0.5B on ECQA
(d) LLaMA-3.1-8B on ECQA
Figure 2: Evolution of predictive information $I(Z_{\ell};Y)$, with lighter curves indicating later training steps. Each curve exhibits a three-phase trend: early layers rise, mid layers peak, and late layers decline.
Figure 2 tracks the trajectory of the predictive information $I(Z_{\ell};Y)$ between the hidden representation at depth $\ell$ and the target label $Y$ throughout training, while Table 1 reports the downstream accuracy obtained when we early exit after a given layer. [^1] The in distribution split is generated with $K_{\text{id}}{=}13$; out of distribution splits use a uniformly–sampled $K_{\text{ood}}\in\{5,\dots,25\}\setminus\{13\}$. Additional results are in Appendix E.
Three-phase information dynamics.
Across models and tasks, predictive information curves exhibit a consistent three-phase pattern:
Progressive Accrual (early layers). In the initial layers, $I(Z_{\ell};Y)$ gradually increases, corresponding to basic feature extraction without substantial task-level comprehension. This aligns with the near-zero accuracy observed in Table 1 for these layers.
Information Peak (intermediate layers). $I(Z_{\ell};Y)$ continues to rise through the mid-to-upper layer, typically peaking before the final few blocks. For the GPT-2 Small on Synthetic Arithmetic dataset, the peak reaches $I\!\approx\!0.9$ in layers 10-11, coinciding with a jump from $\approx 0\%$ to $85\%$ ID accuracy and $59\%$ OOD accuracy. These layers appear to play a critical role in synthesizing abstract features that are essential for generalization.
Table 1: Information Dynamics and Layer-wise Performance (GPT-2-S, Synthetic). OOD performance declines beyond the generalization ridge.
Layer $\mathbf{I(Z;Y)}$ Test Accuracy (%) All In-Dist. Out-Dist. Layer 1 0.0585 0.00 0.00 0.00 Layer 2 0.0673 0.00 0.00 0.00 Layer 3 0.0969 0.00 0.00 0.00 Layer 4 0.1880 0.00 0.00 0.00 Layer 5 0.3038 0.00 0.00 0.00 Layer 6 0.3991 0.00 0.00 0.00 Layer 7 0.5497 4.00 0.10 7.90 Layer 8 0.8067 19.30 0.00 38.60 Layer 9 0.8604 22.35 1.30 43.40 Layer 10 0.9044 57.00 54.30 59.70 Layer 11 0.8650 70.35 85.10 55.60 Layer 12 0.1783 71.95 93.80 50.10
Representational Compression (final layers). Beyond the peak, $I(Z_{\ell};Y)$ decreases, even as in-distribution accuracy approaches $100\%$. The simultaneous drop in OOD accuracy indicates that the final layers tend to memorize training patterns, sacrificing generalization ability.
Notably, this three-phase progression emerges consistently across both GPT-2 Small and GPT-2 Medium, indicating that the observed information dynamics are robust to architectural scale within the same model family.
Generalization Ridge: Memorization-Generalization Trade-off.
The pronounced “information funnel” around intermediate layers reflects a key trade-off between generalization and memorization, which we term the “generalization ridge”. These layers maximize task-relevant information for generalization, while deeper layers increasingly compress and specialize representations, enhancing in-distribution memorization but reducing robustness. This positions intermediate layers as critical control points for managing this trade-off.
Generalization Ridge Emerges Beyond a Depth Threshold
(a) 7 layers: Monotonic Increase
(b) 9 layers: Shallow Peak
(c) Full model: Gen Ridge
Figure 3: Truncating GPT–2 to 7 layers removes the MI peak; 9–layer variants begin to exhibit a shallow peak, and the full 12–layer model shows a pronounced decline.
We empirically vary the depth of GPT-2 Small models fine-tuned on the CLUTRR dataset to examine how model capacity shapes information dynamics, shown in Figure 3. When the model is truncated to 7 layers, $I(Z;Y)$ increases monotonically—indicating insufficient capacity to develop an information peak. Upon increasing the depth to 9 layers, a shallow peak emerges, signifying the threshold at which the model begins to distinguish generalized features from memorized signals. The full 12-layer model exhibits a clear peak followed by a decline, confirming that the generalization ridge emerges only beyond a certain capacity threshold. These results underscore the role of architectural depth in information dynamics and sufficient capacity is required for the generalization ridge to emerge.
Overfitting Adds Memorization in Final Layers
To probe overfitting dynamics, we intentionally fine-tuned the model beyond the optimal point. In Figure 4, we observe that the Predictive Information $I(Z_{\ell};Y)$ rises again in the final layers—departing from the typical compression phase expected at this stage. Before overfitting, the top layers largely behave as pretrained decoders, focusing on surface-level patterns such as token co-occurrence, syntactic templates, or corpus biases, which are not strongly aligned with the task label—hence their initially low $I(Z_{\ell};Y)$. After overfitting, however, the later-layer rise reflects a shift toward memorization, where the model begins to encode superficial shortcuts or redundant label-specific noise rather than useful task information.
Figure 4: $I(Z_{\ell};Y)$ rises in final layers during overfitting (LLaMA, ECQA).
Semantically Important Attention Peaks Where Information Peaks
This discussion aims to provide an intuitive illustration of our generalization ridge hypothesis, highlighting how attention to task-relevant signal tokens shifts across layers (the generalizable information). Specifically, we computed average signal attention across layers, identified the layer with peak signal attention, and compared it to final-layer signal and last-token attention, results are in Table 2. Signal tokens are defined as task-relevant tokens that the model must focus on to solve the task—for the synthetic dataset, these are tokens that appear after the character ‘S’; for CLUTRR, kinship-related entities; and for ECQA, the token corresponding to the correct answer option. This allows us to quantify where in the network attention to semantically important information is concentrated. Our findings show a pattern that supports our hypothesis: (1) Mid-to-late layers peak in signal attention, coinciding with the predictive information ridge (Figure 2), indicating where generalizable representations are strongest. (2) Final layers show reduced attention to signal tokens, suggesting a shift toward memorization to specific data point rather than predictive information abstraction. For example, in Qwen-2.5-0.5B on ECQA, signal attention peaks at Layer 17 (0.2307) but drops to 0.0104 in the final layer, where last-token attention dominates (0.5015).
Table 2: Average attention statistics: (1) average attention scores over all tokens, (2) average attention to signal tokens, (3) the maximum signal attention and its corresponding layer, (4) signal token attention in the final layer and (5) last token attention in the final layer.
Model Dataset Avg. Attn (All) Avg. Attn (Signal) Layer w/ Highest Avg. Signal Final Avg. Signal Final Avg. Last
GPT-2 Small Synthetic 0.0222 0.0409 8 (0.0740) 0.0562 0.0348
GPT-2 Medium CLUTRR 0.0086 0.0086 18 (0.0270) 0.0124 0.1649
Qwen-2.5-0.5B ECQA 0.0225 0.0317 17 (0.2307) 0.0104 0.2169
LLaMA-3.1-8B ECQA 0.0220 0.0394 21 (0.1595) 0.0094 0.5015
*We remove the first token attention score to mitigate attention sink effects.
5.2 Incremental Information Gain: Information Concentrates at Intermediate Layers
To understand how information accumulates across the network, we compute Incremental Information Gain ($I(\Delta Z_{\ell};Y)$)—the mutual information between each residual transition and the target label embedding. As shown in Figure 5, the resulting layer-wise gains reveal that intermediate layers yield the highest information increases. This concentration of information gain further underscores their central role in encoding those task-relevant features that are essential for supporting generalization. The Incremental Information Gain analysis reveals a clear pattern: middle transformer blocks are key to encoding generalizable task-relevant information, thereby forming a generalization ridge. Conversely, later layers contribute little additional predictive signal, and in some cases, actively reduce alignment with the target embeddings. This diminishing contribution in later layers may suggest a shift from general reasoning to memorization of training-specific patterns. This trend underscores a fundamental trade-off in transformer training dynamics. Additional results are in Appendix E.
Figure 5: Middle blocks are key to encoding generalizable information (GPT-2-S, Synthetic).
Residual Scaling Dynamics via Learnable $\beta$ Coefficients
To deepen our understanding of the Generalization Ridge hypothesis, we examine how modulating the contribution of individual transformer blocks affects information flow, and propose a corollary that links layer-wise contribution to generalization performance under distribution shift.
Corollary (Residual Scaling).
Post-peak residual blocks (i.e., layers beyond the generalization ridge) may encode memorized signals. Suppressing these layers’ contribution via learned residual scaling improves out-of-distribution generalization, while amplifying them degrades it.
We introduce a residual scaling mechanism with learnable scalar coefficient parameters, inspired by prior work on adaptive residual modulation (Liu et al., 2019a; Menghani et al., 2024). Transformer architectures inherently employ residual connections to iteratively refine representations. To isolate and quantify the contribution of each transformer block, we scale these residual connections with layer-specific scaling factors $\beta_{\ell}\in\mathbb{R}_{\geq 0}$, with definition below. Each $\beta_{\ell}$ controls the strength of the residual contribution from layer $\ell$, enabling the model to adaptively emphasize or suppress specific blocks:
$z^{(\ell)}=z^{(\ell-1)}+\beta_{\ell}\cdot\text{block}^{(\ell)}(z^{(\ell-1)}),\qquad\beta_{\ell}\in\mathbb{R}.$
Definition (Residual Scaling Coefficient).
$\beta_{\ell}$ is a learnable scalar parameter associated with transformer layer $\ell$, which modulates the contribution of that layer’s residual output to the model’s forward pass.
We freeze model weights and optimize only the residual scaling coefficient parameters $\beta_{\ell}$, which are initialized to 1. These scalars are trained separately on the (a) in-distribution (ID) split and (b) out-of-distribution (OOD) split. Since no other parameters are updated, the learned $\beta_{\ell}$ serve as a direct diagnostic of the extent to which each layer should be amplified or attenuated to suit the data regime, revealing which layers remain stable across regimes and which adapt strongly to distributional shifts.
(a) GPT-2 Small on Synthetic Arithmetic
(b) Qwen on ECQA
Figure 6: Residual scaling coefficients $\beta_{\ell}$ across all layers. ID training emphasizes later layers, while OOD training has lower final layer weight. Curve shows the mean across 5 independent run, and the shaded region denotes 1-$\sigma$ error bar.
As shown in Figure 6, optimizing the residual scaling coefficient parameters on in-distribution data assigns relatively higher residual weights to the final transformer layers, indicating that in-distribution performance benefits from increased reliance on late-layer residual contributions. This observation suggests that these deeper layers capture refinements that are closely tied to the training distribution. In contrast, under out-of-distribution (OOD) settings, the learned coefficients systematically have relative lower contribution of the final layers compared to the ID settings, revealing that improved generalization is associated with downweighting late-stage residual signals. Additional results are in Appendix E. This pattern supports the interpretation that late-layer residuals become increasingly specialized in memorized patterns tied to the training distribution. Together, these findings provide empirical support for the generalization ridge hypothesis, demonstrating that information flow in transformers reflects a trade-off between generalizable signals and memorized, distribution-specific features.
6 Extension to Multi-Token Senario
We further evaluate information flow under multi-token outputs using the CNN/DailyMail summarization dataset (1.0.0, test) (See et al., 2017; Hermann et al., 2015). We evaluate two pretrained base models: Qwen-2.5-0.5B and LLaMA-3.1-8B, generating 50 tokens per sample.
Kernel Extension for Variable-Length Sequences.
To handle multi-token targets, we extend our matrix-based mutual information (MI) estimation to variable-length sequences. Let $Y_{i}=\{y_{i}^{(t)}\}_{t=1}^{T_{i}}$ denote the target sequence for sample $i$. We employ the following kernel to compute sample-to-sample similarity:
$K(Y_{i},Y_{j})=\frac{1}{T_{i}T_{j}}\sum_{t=1}^{T_{i}}\sum_{t^{\prime}=1}^{T_{j}}\kappa\!\left(y_{i}^{(t)},y_{j}^{(t^{\prime})}\right),$
where $\kappa(\cdot,\cdot)$ is a Gaussian kernel. This formulation aggregates all pairwise token similarities between two target sequences into a single scalar, capturing the overall alignment between two token distributions.
We quantify how much information each layer carries about the gold answer by computing $I(Z_{\ell}^{(t)};Y)$ in Figure 7, where $Z_{\ell}$ denotes the hidden state at the last token position at decoding step $t$ and $Y$ represents the full target sequence.
(a) LLaMA on CNN/Daily for 50 generation steps.
(b) Qwen on CNN/Daily for 50 generation steps.
Figure 7: $I(Z_{\ell}^{(t)};Y)$ across layers (x-axis) and generation steps (left color bar), the right y-axis shows mutual information values. Values are plotted every ten steps. Lines denote the mean across five independent runs, with shaded regions indicating $\pm 1$ standard deviation.
Mutual information across generation steps.
We observe several consistent patterns in the mutual information profiles across generation steps. First, the overall magnitude of $I(Z_{\ell}^{(t)};Y)$ increases as generation progresses: later decoding steps tend to exhibit higher mutual information than earlier ones. This reflects the accumulation of information in the hidden representations as the model conditions on an increasingly informative prefix composed of its previously generated tokens. As generation progresses, the hidden state progressively approach the target sequence distribution.
Second, despite differences in magnitude, the layer-wise structure of the mutual information curves remains consistent across generation steps. In particular, $I(Z_{\ell}^{(t)};Y)$ peaks at early-to-intermediate layers and gradually declines toward later layers. The concentration of mutual information in early-to-intermediate layers reflects their role in forming task-relevant representations under autoregressive decoding. These layers integrate information from the input and the evolving context into relatively abstract features that are shared across decoding steps, such as semantic content and task-level structure. In contrast, representations in later layers are increasingly shaped by step-specific decoding requirements, including local token selection, memorized pretrained corpus-specific surface patterns and distribution sharpening. While these representations are critical for producing the next output token, later layers increasingly compress information into a lower-entropy, decision-focused subspace, leading to reduced mutual information with the target sequence.
Figure 8: Mutual information comparing stopwords/punctuation against content words.
Figure 8 compares $I(Z_{\ell};Y)$ when generating content words versus stopwords and punctuation. Both models exhibit similar layer-wise trends across the two token types, with curves largely overlapping throughout early and middle layers, suggesting that intermediate layers perform semantic processing regardless of output token type.
A divergence emerges in the final layers: content words show higher $I(Z_{\ell};Y)$ than stopwords. This reflects differences in lexical selection complexity. Content words draw from a large vocabulary space and require precise semantic discrimination, demanding more target information. Stopwords follow predictable syntactic patterns determined by local grammatical constraints, requiring less target information for selection.
7 Conclusion
We introduce InfoRidge, an information-theoretic framework designed to trace and quantify how information evolves across layers in transformer-based language models for natural language generation. By estimating both predictive information and incremental information gain, we systematically characterize the layerwise dynamics of information flow, offering a principled view of how signals are refined, amplified, or diminished as they propagate through the network. Our findings reveal a consistent generalization ridge emerging in intermediate layers, where mutual information between the hidden representation and the target label reaches its peak before gradually declining. This phenomenon reflects a fundamental trade-off between generalization and memorization as information flows deeper into the model. A similar pattern is observed in the multi-token setting. Attention analysis, capacity analysis and residual scaling experiments offer complementary causal evidence supporting this observation. Taken together, these findings position InfoRidge as a comprehensive framework for diagnosing how language models internally manage information during natural language generation, while also revealing the structural mechanisms that govern the balance between generalization and memorization.
Limitation
Our analysis considers a limited set of datasets and models. In addition, the kernel design for multi-token scenarios may not optimally capture the ground-truth distribution. Future work will extend analysis to broader model families and tasks and explore improved kernel formulations.
Appendix A Mathematical Details for Matrix-Based Information Estimation and Theoretical Foundations
We employ the matrix-based Rényi entropy  Giraldo et al. (2014) to estimate mutual information between representations and labels, leveraging kernel Gram matrices to capture sample similarity.
A.1 Matrix-Based Entropy Estimation
Let $U=\{u_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$ denote $\ell_{2}$-normalized representations obtained from a specific transformer layer, a residual update, or the embedding of the target label. A Gaussian kernel Gram matrix $G_{U}\in\mathbb{R}^{N\times N}$ is constructed as: $(G_{U})_{ij}=\exp\left(-\frac{\|u_{i}-u_{j}\|^{2}}{2\sigma^{2}}\right),$ with bandwidth $\sigma=1$. The matrix is then trace-normalized to ensure $\text{tr}(G_{U})=1$.
The matrix-based Rényi entropy of order $\alpha=1$ is defined as:
$H(U)=-\text{tr}(G_{U}\log G_{U}).$
This expression can be interpreted in terms of the eigenvalue spectrum $\{\lambda_{k}\}$ of $G_{U}$, since $G_{U}$ is positive semi-definite and trace-normalized:
$H(U)=-\sum_{k=1}^{N}\lambda_{k}\log\lambda_{k}.$
The entropy thus reflects the dispersion of the eigenvalues. A more uniform spectrum (i.e., higher entropy) suggests more diversity in the representation space, while a sharply peaked spectrum (i.e., low entropy) indicates redundancy or compression.
Polynomial Kernel. Given a set of normalized representations $U=\{u_{i}\}_{i=1}^{N}$, the polynomial kernel Gram matrix is defined as $(G^{\mathrm{poly}}_{U})_{ij}=\left(u_{i}^{\top}u_{j}+c_{0}\right)^{p},$ where $p$ is the polynomial degree and $c_{0}$ is a constant bias term. Laplacian Kernel. Given the same representation set $U=\{u_{i}\}_{i=1}^{N}$, the Laplacian kernel Gram matrix is defined as $(G^{\mathrm{lap}}_{U})_{ij}=\exp\!\left(-\gamma\,\|u_{i}-u_{j}\|_{1}\right),$ where $\gamma>0$ is a kernel coefficient and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ distance.
A.2 Mutual Information Estimation
To estimate the mutual information between two random variables $U$ and $V$, we compute their Gram matrices $G_{U}$ and $G_{V}$, and form the joint similarity matrix via element-wise (Hadamard) product: $G_{UV}=G_{U}\circ G_{V}.$ After trace-normalization, mutual information is estimated by: I(U;V)=H(U)+H(V)H(U,V),I(U;V)=H(U)+H(V)-H(U,V), where $H(U,V)=-\text{tr}(G_{UV}\log G_{UV})$. The eigenvalue spectrum of $G_{UV}$ governs the joint entropy term; its shape reflects how much of the structure in $U$ and $V$ aligns. A more concentrated spectrum in $G_{UV}$ relative to $G_{U}$ and $G_{V}$ implies stronger dependence and thus higher mutual information.
Appendix B Dataset Overview and Statistics
Data Statistics
We conduct experiments on four datasets with varying levels of complexity and structure: CLUTRR (Sinha et al., 2019), ECQA (Aggarwal et al., 2021), CNN/DailyMail (See et al., 2017; Hermann et al., 2015) and a custom-designed Synthetic Arithmetic dataset. Table 3 summarizes key dataset statistics.
Table 3: Dataset Statistics
Dataset #Train Train Seq. Len #Val Val Seq. Len #Test Test Seq. Len
CLUTRR 9,074 30 2,020 29 1,146 70
ECQA 7,598 21 1,090 21 2,194 21
CNN/DailyMail - - - - 11,490 3,967
Synthetic Arithmetic 10,000 9 2,000 9 2,000 9
B.1 CLUTRR
CLUTRR (Compositional Language Understanding and Text-based Relational Reasoning) (Sinha et al., 2019) is a diagnostic benchmark for evaluating relational inference in language models. Each example contains a story describing family relations, and the task is to infer the missing relationship between two entities. The distribution shift stems from clause lengths that are absent in the training set but present during evaluation. We use the task split “gen_train23_test2to10”, where the model is trained on clause lengths 2 and 3 and evaluated on lengths 2 through 10.
B.2 ECQA
ECQA (Explanations for CommonsenseQA) (Aggarwal et al., 2021) is a commonsense multiple-choice question-answering dataset, where each question is accompanied by 5 answer options.
B.3 CNN/DailyMail
CNN/DailyMail (See et al., 2017; Hermann et al., 2015) is an summarization dataset containing news articles as written by journalists at CNN and the Daily Mail.
B.4 Synthetic Arithmetic Dataset
We construct a synthetic diagnostic dataset to disentangle task-relevant signal learning from spurious noise memorization in a controlled setting. Each sample consists of a sequence of 10 symbolic elements, where the signal component follows an arithmetic progression modulo $K$, and the remainder of each element is independently corrupted with random noise. By varying the modulus $K$, we systematically control task complexity and introduce structured shifts in the data distribution.
Synthetic Arithmetic Dataset Construction.
At each position $t$, the signal value is computed as:
$s_{t}=(s_{0}+t\cdot d)\bmod K,\quad\text{with }s_{0}\in[0,K{-}1],\quad d\in[1,K{-}1].$
Each element in the sequence is represented as a string of the form:
$\texttt{S\{signal\}\_N\{noise\}},\quad\text{where noise}\sim\mathcal{U}_{\text{int}}(0,\texttt{noise\_range-1}).$
Here $\mathcal{U}_{\text{int}}$ denotes the uniform distribution. The model is trained to predict the signal value of the final (10th) element, using the preceding elements as input context.
For example, with $K=5$, $s_{0}=1$, and $d=2$, a sample might look like:
S1_N42 S3_N77 S0_N18 S2_N56 S4_N90 S1_N11 S3_N65 S0_N23 S2_N37
Each element encodes both a signal (the number following S) and a noise component (the number following N). The target is 4, corresponding to the signal of the final (10th) item in the sequence.
Appendix C Implementation Details
This appendix outlines implementation details in our experiments.
C.1 Prompt Construction
All tasks are cast into a natural language generation format. The model receives a prompt and generate the next token. Below are construction strategies and examples for each dataset:
CLUTRR
Each input example in CLUTRR consists of a short narrative describing a set of family relationships, along with a query involving a pair of entities. We construct prompts by concatenating the narrative and a structured natural language question derived from the query tuple. The model is trained to predict the correct relationship as the next token.
Prompt:
Story: [Alice] is [Bob]’s mother. [Bob] is [Charlie]’s father.
Query: What is the relationship between Alice and Charlie? Answer:
Target: grandmother
ECQA (Explanation-augmented Commonsense QA)
Each ECQA instance consists of a multiple-choice question with five candidate answers. We format the prompt by presenting the question followed by all five options (labeled A–E), and conclude with an explicit answer query. The model is trained to predict the correct answer letter as the next token.
Prompt:
Question: What do people usually do at a birthday party?
Options:
A. Sleep
B. Celebrate
C. Cook
D. Exercise
E. Drive
Answer:
Target: B
CNN / DailyMail
[…]
Prompt:
(CNN)The Palestinian Authority officially became the 123rd member of the International Criminal Court on …
The following is a one-paragraph highlight summary of the article:
Target: Membership gives the ICC jurisdiction over alleged crimes committed in Palestinian territories since last June . Israel and the United States opposed the move, which could open the door to war crimes investigations against Israelis .
Synthetic Arithmetic
Each synthetic sample consists of a sequence of 10 symbolic elements, where each element is formatted as S{signal}_N{noise}. The signal values follow an arithmetic progression modulo $K$, and the noise values are independently sampled from a uniform distribution with a fixed range of 100. During training, the modulus $K$ is set to 13. For evaluation, test sequences are generated using values of $K$ from the range $[5,26]$ excluding 13 to simulate a distribution shift. The model receives the first 9 tokens as input and is trained to predict the signal component of the 10th token.
Prompt:
S1_N42 S3_N88 S5_N20 S7_N10 S9_N65 S11_N43 S0_N99 S2_N38 S4_N77
Target: 6
This controlled format enables manipulation of distributional properties by varying the modulus $K$.
C.2 Residual Scaling with Learnable $\beta_{\ell}$ Parameters
We introduce a vector of learnable scalar weights $\beta=\{\beta_{1},\dots,\beta_{L}\}$ applied to residual connections in a frozen transformer:
$z^{(\ell)}=z^{(\ell-1)}+\beta_{\ell}\cdot\text{Block}^{(\ell)}(z^{(\ell-1)}).$
  • All transformer weights are frozen; only $\beta_{\ell}$ parameters are trained.
  • Each transformer block’s residual output is modulated by $\beta_{\ell}$ via forward hooks.
  • $\beta$ are initialized to 1 and updated using gradient descent without altering the architecture.
For Synthetic Arithmetic, the in-distribution (ID) setting is using $K=13$, the out-of-distribution (OOD) setting is using $K=[5,26]$ excluding 13. For CLUTRR, the in-distribution (ID) setting is using the “1.2” and “1.3” split of “gen_train23_test2to10”. Out-of-distribution (OOD) is conducted on held-out CLUTRR configurations that do not overlap with the training distribution: the “1.4” to “1.10” split of “gen_train23_test2to10”. For ECQA, we use training set as the ID setting and CLUTRR as the OOD setting.
C.3 Training and Evaluation Settings
Table 4: Performance (%) of different models across datasets.
Model CLUTRR Synthetic ECQA
GPT-2 Small 56.57 71.95
GPT-2 Medium 62.11 73.15
Qwen-2.5-0.5B 58.61 88.80 65.45
LLaMA-3.1-8B 72.62 89.90 80.49
We finetune the model using HuggingFace’s Trainer API with batch sizes of 16, the learning rate of $\alpha=5\times 10^{-6}$, and AdamW optimizer with weight decay of 0.01. We perform residual $\beta_{\ell}$ training on top of a model that has already been fine-tuned on the target dataset, keeping all previously learned weights frozen and optimizing only the $\beta_{\ell}$ parameters. We use batch sizes of 4, the learning rate of 0.001, Adam optimizer and epochs of 10. Evaluation is conducted by comparing the next predicted token with the next ground-truth target token.
Appendix D Compute and Licensing Details
Computing Resources
Experiments were conducted on an NVIDIA RTX A6000 GPU (48GB).
Model Licenses
The GPT-2 Small (Radford et al., 2019) and GPT-2 Medium (Radford et al., 2019) models are released under the MIT License and are available via Hugging Face Transformers. The Qwen-2.5 0.5B (Yang et al., 2024) model is provided by Alibaba Group under the Apache 2.0 License. The LLaMA-3.1 8B (Meta AI, 2024) model is made available by Meta under a non-commercial research license and accessed via Hugging Face.
Dataset Licenses
CLUTRR (Sinha et al., 2019) is released under a CC BY 4.0 license as part of EMNLP 2019. The ECQA (Aggarwal et al., 2021) dataset is also released under a CC BY 4.0 license by its authors as part of EMNLP 2021. CNN/DailyMail is under Apache-2.0 License. The Synthetic Arithmetic dataset is custom-designed by the authors and does not rely on any external or licensed data sources.
All assets were used in compliance with their respective licenses, and no proprietary or restricted resources were employed in our experiments.
Appendix E Full Quantitative Results with Confidence Estimates
Predictive Information
We subsample 100 test examples per experiment to compute mutual information and report results with statistical significance based on 5 seeds.
(a) GPT-2 Small on CLUTRR
(b) GPT-2 Small on Synthetic Arithmetic
(c) GPT-2 Medium on CLUTRR
(d) GPT-2 Medium on Synthetic Arithmetic
(e) Qwen-2.5-0.5B on ECQA
(f) Qwen-2.5-0.5B on CLUTRR
(g) Qwen-2.5-0.5B on Synthetic Arithmetic
(h) LLaMA-3.1-8B on ECQA
(i) LLaMA-3.1-8B on CLUTRR
(j) LLaMA-3.1-8B on Synthetic Arithmetic
Figure 9: Predictive information $I(Z;Y)$ across models and datasets. Lighter line colors represent later training steps. Each curve shows the mean across 5 random seeds (0, 1, 2, 3, 42), and the shaded region denotes a 2-sigma (~96%) confidence interval.
In cases where the task is too simple relative to model capacity—such as the synthetic arithmetic task with LLaMA—this trend reflects an overfitting regime.
Incremental Information Gain
In addition to predictive information $I(Z^{(\ell)};Y)$, we compute the incremental information gain $I(\Delta Z^{(\ell)};Y)$ for each transformer block.
(a) GPT-2 Small on CLUTRR
(b) GPT-2 Small on Synthetic Arithmetic
(c) GPT-2 Medium on CLUTRR
(d) GPT-2 Medium on Synthetic Arithmetic
(e) Qwen-2.5-0.5B on ECQA
(f) Qwen-2.5-0.5B on CLUTRR
(g) Qwen-2.5-0.5B on Synthetic Arithmetic
(h) LLaMA-3.1-8B on ECQA
(i) LLaMA-3.1-8B on CLUTRR
(j) LLaMA-3.1-8B on Synthetic Arithmetic
Figure 10: Incremental information gain $I(\Delta Z;Y)$ across different models and datasets with ~96% CI error bars.
Kernel Ablation
To assess robustness with respect to kernel selection, we additionally test the Laplacian and Polynomial kernels (Figure 11).
(a) Gaussian Kernel: Qwen-2.5-0.5B on ECQA.
(b) Polynomial Kernel: Qwen-2.5-0.5B on ECQA.
(c) Laplacian Kernel: Qwen-2.5-0.5B on ECQA.
Figure 11: Different kernel shows similar pattern.
Residual Scaling
We present the complete set of residual scaling results, detailing the learned $\beta_{\ell}$ values across all transformer layers.
(a) GPT-2 Small on Synthetic Arithmetic
(b) GPT-2 Medium on CLUTRR
(c) Qwen-2.5-0.5B on ECQA
(d) LLaMA-3.1-8B on ECQA
Figure 12: Residual scaling coefficients $\beta_{\ell}$ across all transformer layers. Each curve shows the mean across five random seeds, and the shaded region denotes 1-sigma error bar. ID training emphasizes later layers, while OOD training has lower final layer weight.